# Conditional success of adaptive therapy: the role of treatment-pausing thresholds revealed by mathematical modeling SUMMARY

### v1

@@ RESULTS
Furthermore, similar to the previous results, considering the uncertainty in real-world scenarios, the probability of adaptive therapy being the best strategy is explored under conditions of random tumor growth rates and status monitoring. The probability that adaptive therapy results in a longer TTP than both MTD and intermittent therapy is shown in FigureΒ 3D1-D2. In real scenarios, since the tumor growth rate is unknown and tumor composition is difficult to obtain, the dependence of TTP on treatment-pausing thresholds, initial tumor burden, and initial resistant cell proportion is explored under random variations of other factors. FigureΒ 3D1 shows the dependence of the probability of adaptive therapy outperforming other treatment strategies on different thresholds ( and initial tumor burden (), with the other parameters, including the initial resistant cell proportion () and the intrinsic growth rates of tumor cells ( and ) varying randomly. FigureΒ 3D2 shows the dependence of the probability on the intial resistant fraction () and the initial tumor burden (), with other parameters, , and , varying randomly.
It can be observed that the probability of selecting adaptive therapy increases significantly when the initial tumor burden is larger. However, when the initial tumor burden is smaller, adaptive therapy does not consistently improve TTP, which aligns with previous findings. Notably, the selection of the treatment-pausing threshold and tumor composition has a weaker impact on the selection of adaptive therapy than the initial tumor burden, indicating that when detailed information about tumor components is unavailable, parameters such as the treatment threshold have less influence, with the initial tumor burden playing a key role. On the other hand, TTP is highly sensitive to the selection of the threshold, among other factors (Figures 2 and S2-S6). In other words, the selection of the optimal threshold is also dependent on various tumor states. Therefore, when other parameters are randomly varied, the impact of the threshold on TTP becomes overshadowed by fluctuations in these parameters.
@@
Conditional superiority landscapes of adaptive therapy compared to other strategies
@@
Under strong competition (, ), untreated tumors evolve toward exclusive dominance of either or , depending on initial composition (FigureΒ 4A), mathematically equivalent to equations (1)β(3) possessing two stable steady states. Continuous MTD administration drives the tumor toward a stable steady state where sensitive and resistant cells maintain a fixed ratio with unchanged tumor burden (FigureΒ 4B). In the absence of treatment, the tumorβs final state depends on its initial condition, mathematically determined by the basin of attraction in which the system starts. Thus, sensitive cell dominance at initiation leads to resistant cell elimination without treatment, requiring only therapeutic containment to prevent long-term progression. Conversely, initial resistant cell dominance under MTD results in stable tumor burden with constant cellular proportions. The initial tumor burden determines whether disease progression occurs. Even when exceeding progression thresholds, tumor burden stabilizes below carrying capacity. This property renders the βuniformly-improveβ scenario inapplicable for strong competition systems, as therapeutic outcomes primarily depend on initial tumor state rather than threshold calibration (FigureΒ 4C). That is, as long as the long-term steady state under treatment remains acceptable (FigureΒ 4B), treatment strategy selection becomes less critical for strong competition cases.
@@ DISCUSSION
Adaptive therapy crucially leverages competitive dynamics to maintain therapeutic-sensitive cell reservoirs that suppress resistant clones, strategically delaying drug resistance evolution while minimizing tumor progression 16. By preserving sensitive cells through threshold-modulated treatment holidays, adaptive therapy sustains ecological suppression of resistant populationsβan advantage over traditional maximum tolerated dose regimens, which inadvertently accelerate resistance 29. The clinical success of adaptive therapy hinges on its paradigm-shifting principle: harnessing intra-tumor competition to transform therapy-sensitive cells into permanent controllers of resistant cell expansion 33, 67. However, the rule-of-thumb thresholds used in clinical practice may require further research and optimization, as they might not fully account for the complex, dynamic interactions between sensitive and resistant cell populations, and may vary across different tumor types and patient profiles 36, 39, 40, 68, 37, 38.
@@
This study employs a Lotka-Volterra model to investigate how treatment-pausing thresholds influence adaptive therapy outcomes in tumors composed of drug-sensitive and resistant cells. By simulating tumor dynamics under varying thresholds, initial tumor burdens, and competition coefficients, the analysis identifies three distinct therapeutic scenarios: uniform-improve (where AT consistently outperforms MTD regardless of threshold selection), conditional-improve (AT efficacy depends on threshold calibration), and uniform-decline (MTD superiority). When considering clinical uncertainties in measurement data, generally, the greater the initial tumor burden, the more likely adaptive therapy will outperform other treatment strategies. However, in cases where the tumor burden is excessively high, the advantages of adaptive therapy become irrelevant, and alternative approaches, beyond pharmacological treatment, may be required.
@@
The competitive dynamics between sensitive and resistant cells is also a key factor in tumor evolution under therapeutic stress. In strong competition environments, tumors tend to evolve toward a state dominated by either sensitive or resistant cells, depending on the initial composition. Continuous MTD therapy drives the tumor toward a stable state where the ratio of sensitive to resistant cells remains constant. Thus, the critical factor in treatment outcomes is not the treatment-pausing threshold but rather the impact of the initial tumor state on the final result. In weak competition environments, sensitive and resistant cells can coexist, and adaptive therapy may lead to resistant cell dominance through Darwinian selection. Treatment effectiveness is influenced by the competition coefficients, and , with lower and higher favoring adaptive therapy. Adaptive therapy prolongs TTP most effectively when sensitive cells maintain a competitive advantage and when initial tumor burden is high, enabling ecological suppression of resistant cells.
@@
This study advances our understanding of adaptive therapy by systematically quantifying the role of treatment-pausing thresholds in modulating tumor evolution under treatment. By integrating ecological competition principles into therapeutic design, the current work provides a mechanistic basis for optimizing intermittent dosing strategies. The findings highlight that adaptive therapy is not universally superior to MTD; its success depends critically on tumor initial state and the treatment-pausing threshold selection. These insights bridge theoretical ecology and clinical oncology, emphasizing the need for dynamic, patient-specific treatment protocols rather than static dosing regimens.
@@
The primary contribution of this work lies in establishing a quantitative relationship between treatment-pausing thresholds and progression-free survival. By categorizing outcomes into three distinct scenarios, the model offers clinicians a potential framework to stratify patients based on measurable parameters such as initial tumor burden and resistant cell fraction. Furthermore, the demonstration that strong competition can stabilize tumor burden indefinitely under specific thresholds has profound implications for managing advanced cancers. These results extend prior work 16, 42, 36, 39 by incorporating threshold-driven mechanism in adaptive therapy.
@@
Despite these advances, the study has limitations. First, the binary classification of tumor cells overlooks phenotypic plasticity and continuous heterogeneity, which are critical drivers of therapeutic resistance 13, 14, 69. Second, model parameters such as competition coefficients (, ) are derived from population-level assumptions rather than molecular data, limiting their biological interpretability 70, 71. Third, the Lotka-Volterra framework assumes strict bistability in competitive interactions, neglecting intermediate scenarios where partial resistance or microenvironmental constraints alter dynamics 72, 73, 74. Future studies should integrate single-cell omics to parameterize competition coefficients mechanistically and explore hybrid models combining continuum and spatial interactions. Clinical validation through adaptive therapy trials incorporating real-time biomarker monitoring will be essential to translate these insights into practice.
@@
Materials availability
This study did not generate new materials.
@@ ACKNOWLEDGMENTS
Conceptualization: Changjing Zhuge, Lei Zhang; methodology: Changjing Zhuge, Haifeng Zhang; investigation, Lanfei Sun, Changjing Zhuge, Kai Kang, Xiao Wang, Haifeng Zhang; writing-βoriginal draft: Lanfei Sun, Changjing Zhuge, Lei Zhang, Haifeng Zhang; writing-βreview & editing: Leyi Zhang, Xiaoxin Wang, Kai Kang, Yanan Cai, Haifeng Zhang, Lei Zhang and Changjing Zhuge; supervision: Changjing Zhuge
@@
The discontinuous changes in Time to Progression (TTP) under neutral competition conditions as a result of threshold variations. The number of treatment pauses has been altered.
- Figures S2-S6:

### v2

@@ RESULTS
Furthermore, similar to the previous results, considering the uncertainty in real-world scenarios, the probability of adaptive therapy being the best strategy is explored under conditions of random tumor growth rates and status monitoring. The probability that adaptive therapy results in a longer TTP than both MTD and intermittent therapy is shown in FigureΒ 3D1-D2. In real scenarios, since the tumor growth rate is unknown and tumor composition is difficult to obtain, the dependence of TTP on treatment-holiday thresholds, initial tumor burden, and initial resistant cell proportion is explored under random variations of other factors. FigureΒ 3D1 shows the dependence of the probability of adaptive therapy outperforming other treatment strategies on different thresholds ( and initial tumor burden (), with the other parameters, including the initial resistant cell proportion () and the intrinsic growth rates of tumor cells ( and ) varying randomly. FigureΒ 3D2 shows the dependence of the probability on the initial resistant fraction () and the initial tumor burden (), with other parameters, , and , varying randomly.
It can be observed that the probability of selecting adaptive therapy increases significantly when the initial tumor burden is larger. However, when the initial tumor burden is smaller, adaptive therapy does not consistently improve TTP, which aligns with previous findings. Notably, the selection of the treatment-holiday threshold and tumor composition has a weaker impact on the selection of adaptive therapy than the initial tumor burden, indicating that when detailed information about tumor components is unavailable, parameters such as the treatment threshold have less influence, with the initial tumor burden playing a key role. On the other hand, TTP is highly sensitive to the selection of the threshold, among other factors (Figures 2 and S3-S7). In other words, the selection of the optimal threshold is also dependent on various tumor states. Therefore, when other parameters are randomly varied, the impact of the threshold on TTP becomes overshadowed by fluctuations in these parameters.
@@
Conditional superiority landscapes of adaptive therapy compared to other strategies under non-neutral competition
@@
Under strong competition (, ), untreated tumors evolve toward exclusive dominance of either or , depending on initial composition (FigureΒ 4A), mathematically equivalent to equations (1)β(3) possessing two stable steady states. Continuous MTD administration drives the tumor toward a stable steady state where sensitive and resistant cells maintain a fixed ratio with unchanged tumor burden (FigureΒ 4B). In the absence of treatment, the tumorβs final state depends on its initial condition, mathematically determined by the basin of attraction in which the system starts. Thus, sensitive cell dominance at initiation leads to resistant cell elimination without treatment, requiring only therapeutic containment to prevent long-term progression. Conversely, initial resistant cell dominance under MTD results in a stable tumor burden with constant cellular proportions. The initial tumor burden determines whether disease progression occurs. Even when exceeding progression thresholds, tumor burden stabilizes below carrying capacity. This property renders the βuniformly-improveβ scenario inapplicable for strong competition systems, as therapeutic outcomes primarily depend on initial tumor state rather than threshold calibration (FigureΒ 4C). That is, as long as the long-term steady state under treatment remains acceptable (FigureΒ 4B), treatment strategy selection becomes less critical for strong competition cases.
@@ DISCUSSION
Adaptive therapy crucially leverages competitive dynamics to maintain therapeutic-sensitive cell reservoirs that suppress resistant clones, strategically delaying drug resistance evolution while minimizing tumor progression 16. By preserving sensitive cells through threshold-modulated treatment holidays, adaptive therapy sustains ecological suppression of resistant populationsβan advantage over traditional maximum tolerated dose regimens, which inadvertently accelerate resistance 29. The clinical success of adaptive therapy hinges on its paradigm-shifting principle: harnessing intra-tumor competition to transform therapy-sensitive cells into permanent controllers of resistant cell expansion 33, 67. However, the rule-of-thumb thresholds used in clinical practice may require further research and optimization as they might not fully account for the complex, dynamic interactions between sensitive and resistant cell populations, and may vary across different tumor types and patient profiles 36, 39, 40, 68, 37, 38.
@@
This study employs a Lotka-Volterra model to investigate how treatment-holiday thresholds influence adaptive therapy outcomes in tumors composed of drug-sensitive and resistant cells. By simulating tumor dynamics under varying thresholds, initial tumor burdens, and competition coefficients, the analysis identifies three distinct therapeutic scenarios: uniform-improve (where AT consistently outperforms MTD regardless of threshold selection), conditional-improve (AT efficacy depends on threshold calibration), and uniform-decline (MTD superiority). When considering clinical uncertainties in measurement data, generally, the greater the initial tumor burden, the more likely adaptive therapy will outperform other treatment strategies. However, in cases where the tumor burden is excessively high, the advantages of adaptive therapy become irrelevant, and alternative approaches, beyond pharmacological treatment, may be required.
@@
The competitive dynamics between sensitive and resistant cells is also a key factor in tumor evolution under therapeutic stress. In strong competition environments, tumors tend to evolve toward a state dominated by either sensitive or resistant cells, depending on the initial composition. Continuous MTD therapy drives the tumor toward a stable state where the ratio of sensitive to resistant cells remains constant. Thus, the critical factor in treatment outcomes is not the treatment-holiday threshold but rather the impact of the initial tumor state on the final result. In weak competition environments, sensitive and resistant cells can coexist, and adaptive therapy may lead to resistant cell dominance through Darwinian selection. Treatment effectiveness is influenced by the competition coefficients, and , with lower and higher favoring adaptive therapy. Adaptive therapy prolongs TTP most effectively when sensitive cells maintain a competitive advantage and when initial tumor burden is high, enabling ecological suppression of resistant cells.
@@
This study advances our understanding of adaptive therapy by systematically quantifying the role of treatment-holiday thresholds in modulating tumor evolution under treatment. By integrating ecological competition principles into therapeutic design, the current work provides a mechanistic basis for optimizing intermittent dosing strategies. The findings highlight that adaptive therapy is not universally superior to MTD; its success depends critically on tumor initial state and the treatment-holiday threshold selection. These insights bridge theoretical ecology and clinical oncology, emphasizing the need for dynamic, patient-specific treatment protocols rather than static dosing regimens.
@@
This work establishes a quantitative relationship between treatment-holiday thresholds and progression-free survival. By categorizing outcomes into three distinct scenarios, the model offers clinicians a potential framework to stratify patients based on measurable parameters such as initial tumor burden and resistant cell fraction. Furthermore, the demonstration that strong competition can stabilize tumor burden indefinitely under specific thresholds has profound implications for managing advanced cancers. These results extend prior work 16, 42, 36, 39 by incorporating threshold-driven mechanism in adaptive therapy.
@@
Despite these advances, the study has several limitations. First, the binary classification of tumor cells overlooks the phenotypic plasticity and continuous heterogeneity that are critical drivers of therapeutic resistance 13, 14, 69. Second, model parameters such as the competition coefficients (, ) are based on population-level assumptions rather than molecular data, necessitating further investigation into the microscopic foundations of these assumptions 70, 71. Third, the Lotka-Volterra model does not account for the spatial structure of tumors 72 or the microenvironmental constraints that alter tumor dynamics 73, 74. Future studies should integrate single-cell and spatial omics data to mechanistically parameterize competition coefficients and explore hybrid models that combine continuum and spatial interactions. Additionally, clinical validation through adaptive therapy trials, incorporating real-time biomarker monitoring, will be crucial for translating these insights into practice.
@@ ACKNOWLEDGMENTS
Conceptualization: Changjing Zhuge, Lei Zhang; methodology: Changjing Zhuge, Haifeng Zhang; investigation, Lanfei Sun, Changjing Zhuge, Kai Kang, Xiaoxin Wang, Haifeng Zhang; writingβoriginal draft: Lanfei Sun, Changjing Zhuge, Lei Zhang, Haifeng Zhang; writingβreview & editing: Leyi Zhang, Xiaoxin Wang, Kai Kang, Yanan Cai, Haifeng Zhang, Lei Zhang and Changjing Zhuge; supervision: Changjing Zhuge
@@
The discontinuous changes in time to progression (TTP) under neutral competition conditions as a result of threshold variations. The number of treatment holidays has been altered.
- Figures S2-S7: